If normal at one extremity of latus rectum of an ellipse passes through one extremity of minor axis, prove that the eccentricity e is given by e^4+e^2,-1=0
1
Expert's answer
2020-06-21T14:43:58-0400

Let the equation of the ellipse is:

"\dfrac{x^2}{a} + \dfrac{y^2}{b}=1"


The equation of the normal to this ellipse at "\left(ae , \dfrac{b^2}{a}\right)" is:



"\dfrac{x-ae}{\frac{ae}{a^2}} = \dfrac{y-\frac{b^2}{a}}{\frac{b^2}{ab^2}} \ \rightarrow (1) \\\\"



This normal passes through "(0,\u2212b)"  according to given condition.


Substitute "x=0 , y=-b" in equation (1)

"\dfrac{0-ae}{\frac{ae}{a^2}} = \dfrac{-b-\frac{b^2}{a}}{\frac{b^2}{ab^2}}"

Simplify

"-a^2=-ab-b^2 \ \rightarrow (2)"

Since "e^2=1-\dfrac{b^2}{a^2}"


So, "\dfrac{b^2}{a^2}= 1-e^2"


Therefore, "b^2=a^2(1-e^2)"


Substitute "b^2" in equation (2)



"-a^2=-ab-a^2(1-e^2)"

Simplify

"a^2=ab+a^2-a^2e^2"


"ab=a^2e^2"


"b=ae^2"

So,

"b^2 = a^2e^4"

And we know that

"b^2=a^2(1-e^2)"

Therefore,

"a^2e^4=a^2(1-e^2)"

Simplify



"e^4+e^2-1=0"